# Colouring ($P_{2}\cup P_{4}$, diamond)-free graphs with $\omega$ colours

### v1

@@ 1.1 Our Contributions
Through a different structural decomposition approach, we obtain the optimal $\chi$-binding function across all values of $\omega(G)$, thereby providing a complete solution to this colouring problem.
@@ 2 The Structure of ($P_{2}\cup P_{4}$, diamond)-Free Graphs
Since $G$ is diamond-free, each $S_{i}^{j}$ is stable. Note that $\omega(R_{j}),\omega(T_{i})\leq\omega-1$ and $\omega(Z)\leq k$. Furthermore, by Lemma 2.1, each of the sets $R_{j}$, $T_{i}$, and $Z$ induces a perfect graph. We then define
@@
Let $B$ be the $K_{k}$ that is anticomplete to $A$. We first list some properties of $Z$, $R$, $S$, and $T$. The proofs of these properties are analogous to those given in the proof of Claim 2.1, and hence are omitted.
@@ 4 The Case $k\geq 4$
By symmetry, we may assume that $x,y\in(C_{0}\cup C_{1})\cap(C_{0}^{\prime}\cup C_{1}^{\prime})$. Thus, $\{x,y\}$ is anticomplete to $\{a_{2},\ldots,a_{\omega},b_{2},\ldots,b_{k}\}$. By ($\ast$2), $\{a_{2},\ldots,a_{\omega}\},\{b_{2},\ldots,b_{k}\}]$ is a matching. Since $\omega,k\geq 4$ and $G$ is $(P_{2}\cup P_{4})$-free, Lemma 1.4 implies that $\{a_{2},\ldots,a_{\omega}\}$ is anticomplete to $\{b_{2},\ldots,b_{k}\}$. Since $N_{A\cup B}(x)\neq N_{A\cup B}(y)$, $N_{A}(x)\neq N_{A}(y)$ or $N_{B}(x)\neq N_{B}(y)$. Suppose $N_{A}(x)\neq N_{A}(y)$ (the case where $N_{B}(x)\neq N_{B}(y)$ is analogous). Then either $a_{1}\sim x$ and $a_{1}\nsim y$, or $a_{1}\nsim x$ and $a_{1}\sim y$. By symmetry, assume that $a_{1}\sim x$ and $a_{1}\nsim y$. Since $k\geq 4$, there exist vertices $b_{i},b_{j}\in\{b_{2},\ldots,b_{k}\}$ that are anticomplete to $a_{1}$. Consequently, $\{b_{i},b_{j},a_{2},a_{1},x,y\}$ induces a $P_{2}\cup P_{4}$, a contradiction.
@@
(P6) If $\omega=4$, then the vertex set of each component in $G[S]$ can be partitioned into at most four stable sets, each contained in a distinct $S_{i}^{j}$, with all $i$’s and all $j$’s being distinct within that component.
@@
Let $G$ be a $(P_{2}\cup P_{4},\text{ diamond})$-free graph with $\omega\geq 4$ and $k\geq 4$. By ($\ast\ast$2), $A$ is not anticomplete to $B$. Then, by ($\ast$2), we may relabel $B$ so that, for some integer $m\in[k]$, we have $a_{i}\sim b_{i}$ for all $i\in[m]$, and $a_{p}\nsim b_{q}$ whenever $p\in[\omega]\setminus[m]$ and $q\in[k]\setminus[m]$. By Lemma 2.1, each of the sets $R_{j}$, $T_{i}$, and $Z$ induces a perfect graph. Since $\omega(R_{j}),\omega(T_{i})\leq\omega-1$ and $\omega(Z)\leq k\leq\omega$, we may colour them with at most $\omega-1$, $\omega-1$, and $\omega$ colours, respectively. Now we construct an $\omega$-colouring of $G$.
@@
  • If $\omega\geq 5$, by (P4) and (P7), we may colour each $S_{i}^{j}$ with one colour from $\omega]\setminus(c(A_{i})\cup c(B_{j}))$ for $i\in[\omega]$ and $j\in[k]$.
@@
Data availability
No data was used for the research described in the article.

### v2

@@ 1.1 Our Contributions
Through a different structural decomposition approach, we obtain the optimal $\chi$-binding function across all values of $\omega(G)$, thereby providing a complete solution for the $\chi$-binding function of this class.
@@ 2 The Structure of ($P_{2}\cup P_{4}$, diamond)-Free Graphs
Note that $\omega(R_{j}),\omega(T_{i})\leq\omega-1$ and $\omega(Z)\leq k$. Furthermore, by Lemma 2.1, each of the sets $R_{j}$, $S_{i}^{j}$, $T_{i}$, and $Z$ induces a perfect graph. We then define
@@
Let $B$ be the $K_{k}$ that is anticomplete to $A$. Since $G$ is diamond-free, every $S_{i}^{j}$ is stable. We first list some properties of $Z$, $R$, $S$, and $T$. The proofs of these properties are analogous to those given in the proof of Claim 2.1, and hence are omitted.
@@ 4 The Case $k\geq 4$
By symmetry, we may assume that $x,y\in(C_{0}\cup C_{1})\cap(C_{0}^{\prime}\cup C_{1}^{\prime})$. Thus, $\{x,y\}$ is anticomplete to $\{a_{2},\ldots,a_{\omega},b_{2},\ldots,b_{k}\}$. By ($\ast$2), $\{a_{2},\ldots,a_{\omega}\},\{b_{2},\ldots,b_{k}\}]$ is a matching. Since $\omega,k\geq 4$ and $G$ is $(P_{2}\cup P_{4})$-free, Lemma 1.4 implies that $\{a_{2},\ldots,a_{\omega}\}$ is anticomplete to $\{b_{2},\ldots,b_{k}\}$. Since $N_{A\cup B}(x)\neq N_{A\cup B}(y)$, it follows that $N_{A}(x)\neq N_{A}(y)$ or $N_{B}(x)\neq N_{B}(y)$. Suppose $N_{A}(x)\neq N_{A}(y)$ (the case where $N_{B}(x)\neq N_{B}(y)$ is analogous). Then either $a_{1}\sim x$ and $a_{1}\nsim y$, or $a_{1}\nsim x$ and $a_{1}\sim y$. By symmetry, assume that $a_{1}\sim x$ and $a_{1}\nsim y$. Since $k\geq 4$, there exist vertices $b_{i},b_{j}\in\{b_{2},\ldots,b_{k}\}$ that are anticomplete to $a_{1}$. Consequently, $\{b_{i},b_{j},a_{2},a_{1},x,y\}$ induces a $P_{2}\cup P_{4}$, a contradiction.
@@
(P6) If $\omega=4$ and $S_{i}^{p}$ is stable for all $i\in[\omega]$ and $p\in[k]$, then the vertex set of each component in $G[S]$ can be partitioned into at most four stable sets, each contained in a distinct $S_{i}^{j}$, with all $i$’s and all $j$’s being distinct within that component.
@@
(P8) If there exists an $S_{i}^{p}$ that is not stable, then $A,B]=\{a_{i}b_{p}\}$, and every set $S_{j}^{q}$ with $(j,q)\neq(i,p)$ is stable.
Proof.
By symmetry, it suffices to show that (P8) holds in the case that $S_{1}^{1}$ is not stable. Assume that there exist two adjacent vertices $x,y\in S_{1}^{1}$. Then $a_{1}\sim b_{1}$, since $G$ is diamond-free. By ($\ast$2), $\{a_{2},\ldots,a_{\omega}\},\{b_{2},\ldots,b_{k}\}]$ is a matching. Since $\omega,k\geq 4$ and $G$ is $(P_{2}\cup P_{4})$-free, Lemma 1.4 implies that $\{a_{2},\ldots,a_{\omega}\}$ is anticomplete to $\{b_{2},\ldots,b_{k}\}$. Hence, $A,B]=\{a_{1}b_{1}\}$. Since $G$ is diamond-free, every set $S_{j}^{q}$ with $(j,q)\neq(1,1)$ is stable. ∎
(P9) If there exists an $S_{i}^{p}$ that is not stable, then $S_{j}^{q}$ is anticomplete to $S_{j^{\prime}}^{q^{\prime}}$ for distinct $q,q^{\prime}\in[k]$ and distinct $j,j^{\prime}\in[\omega]$.
Proof.
If $\omega\geq 5$, then (P9) follows directly from (P7). We therefore consider the case $\omega=4$ and $k=4$. By symmetry, assume that $i=p=1$. By (P8), we have $A,B]=\{a_{1}b_{1}\}$. Suppose, to the contrary, that there exist vertices $x\in S_{j}^{q}$ and $y\in S_{j^{\prime}}^{q^{\prime}}$ such that $x\sim y$, where $j\neq j^{\prime}$ and $q\neq q^{\prime}$. Since $A,B]=\{a_{1}b_{1}\}$ and $\omega=k=4$, we have
$|[A\setminus\{a_{j},a_{j^{\prime}}\},\,B\setminus\{b_{q},b_{q^{\prime}}\}]|\leq 1.$
Thus, there exists a vertex $w\in A\setminus\{a_{j},a_{j^{\prime}}\}$, with $w\neq a_{1}$, that is anticomplete to $B\setminus\{b_{q},b_{q^{\prime}}\}$. If $a_{j}\neq a_{1}$, then $\{w,a_{j},x,y\}\cup(B\setminus\{b_{q},b_{q^{\prime}}\})$ induces a $P_{2}\cup P_{4}$. If $a_{j^{\prime}}\neq a_{1}$, then $\{w,a_{j^{\prime}},x,y\}\cup(B\setminus\{b_{q},b_{q^{\prime}}\})$ induces a $P_{2}\cup P_{4}$. In either case, we obtain a contradiction. ∎
@@
Let $G$ be a $(P_{2}\cup P_{4},\text{ diamond})$-free graph with $\omega\geq 4$ and $k\geq 4$. By ($\ast\ast$2), $A$ is not anticomplete to $B$. Then, by ($\ast$2), we may relabel $B$ so that, for some integer $m\in[k]$, we have $a_{i}\sim b_{i}$ for all $i\in[m]$, and $a_{p}\nsim b_{q}$ whenever $p\in[\omega]\setminus[m]$ and $q\in[k]\setminus[m]$. By Lemma 2.1, each of the sets $R_{j}$, $T_{i}$, and $Z$ induces a perfect graph. Since $\omega(R_{j}),\omega(T_{i})\leq\omega-1$ and $\omega(Z)\leq k\leq\omega$, we may colour them with at most $\omega-1$, $\omega-1$, and $\omega$ colours, respectively. We now construct an $\omega$-colouring of $G$ by distinguishing two cases according to whether every $S_{i}^{p}$ is stable.
Case 1. $S_{i}^{p}$ is stable for all $i\in[\omega]$ and $p\in[k]$.
@@
  • If $\omega\geq 5$, by (P4) and (P7), we may colour each $S_{i}^{j}$ with at most one colour from $\omega]\setminus(c(a_{i})\cup c(b_{j}))$ for $i\in[\omega]$ and $j\in[k]$.
@@
  • By (P1), colour $Z$ with at most $\omega$ colours from $\omega]$.
Case 2. There exists an $S_{i}^{p}$ that is not stable.
By (P8) together with the argument given in the first paragraph of the proof, it follows that $i=p=m=1$, $A,B]=\{a_{1}b_{1}\}$, and every set $S_{j}^{q}$ with $(j,q)\neq(1,1)$ is stable. Note that $\omega(S_{1}^{1})\leq\omega-2$ and $S_{1}^{1}$ is perfect, we may colour it with at most $\omega-2$ colours.
  • Colour $a_{i}$ with colour $i$ for $i\in[\omega]$.
  • Colour $b_{i}$ with colour $i+1$ for $i\in[k-1]$, and colour $b_{k}$ with colour $1$.
  • By (P4) and (P9), colour $S_{1}^{1}$ with at most $\omega-2$ colours from $\omega]\setminus\{1,2\}$, and colour $S_{j}^{q}$ with at most one colour from $\omega]\setminus(c(a_{i})\cup c(b_{j}))$ for $(j,q)\neq(1,1)$.